40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 194     Entry time: Wed Jun 30 23:26:27 2010
Author: tarac 
Type: DailyProgress 
Category: RefCav 
Subject: Debugging FSS servo+Locking RefCav 

summary of this entry

  RefCav was not locked becase of the RFPD's input power. I haven't checked yet if it's the power supply or the cable. Now RefCav is locked.

 

Details

I was trying to lock RefCav again, but it didn't work.

So, I checked the error signal from mixer out, saw nothing.

It turned out that RF out from the PD has low voltage output, ~10mV.

I unplugged the cable and switch to another PD's cable along with its power supply, the DC increased to ~300 mV, a good sign.

Still, no error signal from Mixer Out Channel, nothing at all.

The control loop is not closed. Now 35.5 MHz LO is connected to the EOM, and to the servo card, LO input channel. The RF signal goes to the PD input.

We should be able to see the error signal from this setup.

**********The loop in the medm screen must be enabled in order to see the error signal******.

 

     EOM  -----------------BS--------------------- [ RefCav ]

        |                           |

        LO                       PD

        |                             |

        -------- X -------------

                   |

                   --->  Error signal

 

  I mix the signals from PD and LO by a Mixer [Minicurcuit ZFM 3 S+] and see something. It's a peak crossing zero at center instead of three peaks like the error signals. 

So I'm not sure what I'm seeing, but it seems that Mixer Out channel from the spare FSS servo might not working. I switched back to the first card, and now I see the error signals.

I set the RF to 6.1 V which corresponds to 97 mV of the error signal's P-P height.

Phase adj: is 4.5122 V,

phase flip: 0

Gain: [not set yet]

      All these values are saved in startup

  To summary, RefCav was not locked becase of the power cable which I haven't checked yet if it's the power supply or the cable. Now it's locked.

I'll set the gain tomorrow and see where I can connect the slow actuator for the laser driver. Right now I can't find where the servo for slow actuator would be.

 

ELOG V3.1.3-